# Prosoluble subgroups of the profinite completion of the fundamental group of compact 33 33-manifolds

### v2

@@ 8.3 Geometric cases
The Geometrization conjecture, proved by G. Perelman on his approach to prove the Poincaré conjecture (see [Per02], [Per03b], [Per03a]), says that a 33 33-manifold is geometrically modeled by one of the eight Thurston’s geometries
@@
    a finite cyclic group or a central extension of $D_{2n}$, $T$, $O$ or $I$ by a cyclic group of even order. We are denoting by $T$ the tetrahedral group of 1212 1212 rotational symmetries of a tetrahedron, by $O$ the octahedral group of 2424 2424 rotational symmetries of a cube or of an octahedron, by $I$ the icosahedral group of 6060 6060 rotational symmetries of a dodecahedron or an icosahedron;
@@
    virtually soluble.
The fundamental groups of these first five geometries are soluble. If we consider $H$ as a prosoluble subgroup of the profinite completion of these groups, then $H$ can be any of its subgroups. We summarize the existing classification as:
@@
    $H$ is a subgroup of the profinite completion of a 33 33-dimensional Bieberbach group (i.e, virtually $\mathbb{Z}^{3}$);
@@
    $H$ is a subgroup of the profinite completion of a group extension of $\mathbb{Z}^{2}\rtimes_{A}\mathbb{Z}$ by $C$ where $C$ is trivial or isomorphic to $C_{2}$ and $A$ is an Anosov matrix;
@@ 8.4 Seifert manifolds
In order to conclude the analysis of the possible irreducible compact 33 33-manifolds we now deal with the Seifert manifolds. We have the following exact sequence:
@@
    $H$ is a subgroup of a finite free prosoluble product of finite cyclic $p$-groups;
@@
Since $\widehat{C}$ is malnormal in $\widehat{\pi_{1}(O)}$, by [WZ17, Lemma 7.3], $\widehat{\pi_{1}(O)}$ is a 11 11-acylindrical finite graph of profinite groups. Applying Theorem F we get one of the following cases:
@@
    $H$ is subgroup of the free prosoluble product of $H\cap\widehat{F}^{g}$ or $H\cap\widehat{K}^{g}$ and these intersection are pro-$p$.
@@
Suppose that (a) holds. Note that $H\cap\widehat{F}^{g}$ is free pro-$p$ and $H\cap\widehat{K}^{g}$ is free pro-$p$ product of cyclic pro-$p$ groups by the pro-$p$ version of the Kurosh Subgroup Theorem (cf. [Rib17, Theorem 9.6.2]). Then we have $(i)$. If $H$ is Frobenius, we have item $(ii)$.
@@
Let $G$ be the fundamental group of a Seifert manifold with either geometry $\mathbb{H}^{2}\times\mathbb{R}$ or $\widetilde{\operatorname{SL}_{2}(\mathbb{R})}$ and $H$ a finitely generated prosoluble subgroup of $\widehat{G}$. Then $H$ is an extension of a torsion-free procyclic group ^$\widehat{\mathbb{Z}}_{\sigma}$ by a subgroup of a finite free prosoluble product of finite cyclic $p$-groups.
@@
where $\pi_{1}(O)$ is a Fuchsian group acting on $\mathbb{Z}$ either trivially or by inversion. Note that we have an action of $\widehat{\pi_{1}(O)}$ therefore of $H_{0}=H\widehat{\mathbb{Z}}/\widehat{\mathbb{Z}}$ on $\widehat{\mathbb{Z}}$ induced by the action of $\pi_{1}(O)$ on $\mathbb{Z}$ which is trivial or by inversion.
@@
where ^$\widehat{\mathbb{Z}}_{\sigma}=H\cap\widehat{\mathbb{Z}}$ and $H_{0}$ acts on ^$\widehat{\mathbb{Z}}_{\sigma}$ either trivially or by inversion.
@@
    $H_{0}$ is a subgroup of a finite free prosoluble product of finite cyclic $p$-groups;
@@
    In this case $H$ is an extension of ^$\widehat{\mathbb{Z}}_{\sigma}$ by $H_{0}$ with the either trivial action or by inversion.
@@
    Suppose that
    $1\to\widehat{\mathbb{Z}}\to H\to\widehat{\mathbb{Z}}_{\pi}\rtimes C_{n}\to 1$
    is an exact sequence with ^$\widehat{\mathbb{Z}}_{\pi}\rtimes C_{n}$ a profinite Frobenius group with $C_{n}$ a finite cyclic group of order $n$. As the lift of every element of finite order of $\pi_{1}(O)$ to $\pi_{1}(M)$ must centralize $\mathbb{Z}$, the lift of every element of finite order of $H_{0}$ must centralize ^$\widehat{\mathbb{Z}}_{\sigma}$. But $H_{0}$ is generated by torsion elements, so the extension is central in this case.
    Let $\rho$ be the set of all primes dividing $n$. Then the preimage of $C_{n}$ in $H$ is ^$\widehat{\mathbb{Z}}_{\rho^{\prime}}\times\widehat{\mathbb{Z}}_{\rho}$ where $\rho^{\prime}$ is the complement of $\rho$ in $\sigma$ (it is isomorphic but not equal to ^$\widehat{\mathbb{Z}}_{\sigma}$. Thus we can write $H$ as (^$(\widehat{\mathbb{Z}}_{\rho^{\prime}}\times\widehat{\mathbb{Z}}_{\pi})\rtimes\widehat{\mathbb{Z}}_{\rho}\cong\widehat{\mathbb{Z}}_{\rho^{\prime}}\rtimes(\widehat{\mathbb{Z}}_{\pi}\rtimes\widehat{\mathbb{Z}}_{\rho})$. As $\pi\cap\rho=\emptyset$, ^$\widehat{\mathbb{Z}}_{\pi}\rtimes\widehat{\mathbb{Z}}_{\rho}$ is projective and hence a subgroup of a free profinite product of cyclic $p$-groups.
@@
    Suppose that
    $1\to\widehat{\mathbb{Z}}_{\sigma}\to H\to C_{n}\to 1$
    is an exact sequence with $C_{n}$ finite cyclic of order $n$. As the lift of every element of finite order of $\pi_{1}(O)$ to $\pi_{1}(M)$ must centralize $\mathbb{Z}$, the lift of every element of finite order of $H_{0}$ must centralize ^$\widehat{\mathbb{Z}}_{\sigma}$. Hence the extension is central and so $H$ is procyclic torsion-free.

### v3

@@ 8.3 Geometric cases
The Geometrization conjecture, proved by G. Perelman on his approach to prove the Poincaré conjecture (see [Per02], [Per03b], [Per03a]), says that a $3$-manifold is geometrically modeled by one of the eight Thurston’s geometries
@@
    a finite cyclic group or a central extension of $D_{2n}$, $T$, $O$ or $I$ by a cyclic group of even order. We are denoting by $T$ the tetrahedral group of $12$ rotational symmetries of a tetrahedron, by $O$ the octahedral group of $24$ rotational symmetries of a cube or of an octahedron, by $I$ the icosahedral group of $60$ rotational symmetries of a dodecahedron or an icosahedron;
@@
    soluble.
The fundamental groups in the case (i) are soluble. except the case of an extension of $I$ by a cyclic group of even order. The icosahedral group $I$ is isomorphic to $A_{5}$, which is non-soluble. Recall that $A_{5}$ has no proper subgroups of order greater than $12$ so the possible subgroups are $C_{2},C_{3}$, $C_{5}$, $D_{3},D_{5},A_{4}$ and they are covered in the previous cases. Thus, to analyze the soluble subgroups when $\Gamma$ is a central extension of $A_{5}$ by $C_{2n}$ we can reduce it to the other soluble cases.
If we consider $H$ as a prosoluble subgroup of the profinite completion of groups (i)-(v), then $H$ can be any of its subgroups. We summarize the existing classification as:
@@
    $H$ is a subgroup of the profinite completion of a $3$-dimensional Bieberbach group (i.e, virtually $\mathbb{Z}^{3}$);
@@
    $H$ is a subgroup of the profinite completion of a group extension of $\mathbb{Z}^{2}\rtimes_{A}\mathbb{Z}$ by $C$ where $C$ is trivial or isomorphic to $C_{2}$ and $A\in\operatorname{GL}_{2}(\mathbb{Z})$ or a central extension of $\mathbb{Z}$ by $\mathbb{Z}^{2}$;
@@ 8.4 Seifert manifolds
In order to conclude the analysis of the possible irreducible compact $3$-manifolds we now deal with the Seifert manifolds. We have the following exact sequence:
@@
    $H$ is a subgroup of a free prosoluble product of finitely many finite cyclic $p$-groups;
@@
Since $\widehat{C}$ is malnormal in $\widehat{\pi_{1}(O)}$, by [WZ17, Lemma 7.3], $\widehat{\pi_{1}(O)}$ is a $1$-acylindrical finite graph of profinite groups. Applying Theorem F we get one of the following cases:
@@
    $H$ is a subgroup of the free prosoluble product of intersections $H\cap\widehat{F}^{h}$, $h\in\widehat{\pi_{1}(O)}$ and $H\cap\widehat{K}^{g}$, $g\in\widehat{\pi_{1}(O)}$, and these intersections are pro-$p$.
@@
Suppose that (a) holds. Note that $H\cap\widehat{F}^{g}$ is free pro-$p$ and $H\cap\widehat{K}^{g}$ is free $p pro$duct of cyclic pro-$p$ groups by the pro-$p$ version of the Kurosh Subgroup Theorem (cf. [Rib17, Theorem 9.6.2]). Then we have $(i)$. If $H$ is Frobenius, we have item $(ii)$.
@@
Let $G$ be the fundamental group of a Seifert manifold with either geometry $\mathbb{H}^{2}\times\mathbb{R}$ or $\widetilde{\operatorname{SL}_{2}(\mathbb{R})}$ and $H$ a finitely generated prosoluble subgroup of $\widehat{G}$. Then $H$ is an extension of a torsion-free procyclic group $\widehat{\mathbb{Z}}_{\sigma}$ by a subgroup of a free prosoluble product of finitely many finite cyclic $p$-groups.
@@
where $\pi_{1}(O)$ is a Fuchsian group acting on $\mathbb{Z}$ either trivially or by inversion. Then we have an action of $\widehat{\pi_{1}(O)}$ and hence of $H_{0}=H\widehat{\mathbb{Z}}$$\widehat{\mathbb{Z}}$ on $\widehat{\mathbb{Z}}$ and this action is trivial or by inversion.
@@
where $\widehat{\mathbb{Z}}_{\sigma}=H\cap\widehat{\mathbb{Z}}$ and $H_{0}$ acts on $\widehat{\mathbb{Z}}_{\sigma}$ either trivially or by inversion.
@@
    $H_{0}$ is a subgroup of a free prosoluble product of finitely many finite cyclic $p$-groups;
@@
    In this case $H$ is an extension of $\widehat{\mathbb{Z}}_{\sigma}$ by $H_{0}$ with the either trivial action or by inversion.
@@
    Suppose that 1 →^Z_σ →H →^Z_π ⋊C_n →1 is an exact sequence with $\widehat{\mathbb{Z}}_{\pi}\rtimes C_{n}$ a profinite Frobenius group with $C_{n}$ a finite cyclic group of order $n$. As the lift of every element of finite order of $\pi_{1}(O)$ to $\pi_{1}(M)$ must centralize $\mathbb{Z}$, the lift of every element of finite order of $H_{0}$ must centralize $\widehat{\mathbb{Z}}_{\sigma}$. But $H_{0}$ is generated by torsion elements, so the extension is central in this case.
    Let $\rho$ be the set of all primes dividing $n$. Then the preimage of $C_{n}$ in $H$ is $\widehat{\mathbb{Z}}_{\rho^{\prime}}\times\widehat{\mathbb{Z}}_{\rho}$ where $\rho^{\prime}$ is the complement of $\rho$ in $\sigma$ (it is isomorphic but not equal to $\widehat{\mathbb{Z}}_{\sigma}$). Thus we can write $H$ as ($\widehat{\mathbb{Z}}_{\rho^{\prime}}\times\widehat{\mathbb{Z}}_{\pi})\rtimes\widehat{\mathbb{Z}}_{\rho}\cong\widehat{\mathbb{Z}}_{\rho^{\prime}}\rtimes(\widehat{\mathbb{Z}}_{\pi}\rtimes\widehat{\mathbb{Z}}_{\rho})$). As $\pi\cap\rho=\emptyset$, $\widehat{\mathbb{Z}}_{\pi}\rtimes\widehat{\mathbb{Z}}_{\rho}$ is projective and hence a subgroup of a free profinite product of cyclic $p$-groups (since a finitely generated projective group is a subgroup of a free group of finite rank).
@@
    Suppose that 1 →^Z_σ →H →C_n →1 is an exact sequence with $C_{n}$ finite cyclic of order $n$. As the lift of every element of finite order of $\pi_{1}(O)$ to $\pi_{1}(M)$ must centralize $\mathbb{Z}$, the lift of every element of finite order of $H_{0}$ must centralize $\widehat{\mathbb{Z}}_{\sigma}$. Hence the extension is central and so $H$ is procyclic torsion-free.